I don't think this is a proportion questions, but didn't know what to call it. I am trying to help my 12 year old son.
Rick earns $500 less than three times as much as Jim. If their combined salary is $49,500, how much do they each earn?

Answer by rfer(16322)

jim=x=$12,500
rick=3x-500=$37,000
----------------------
x+3x-500=49500
add +500 to both sides
x+3x=50000
combine like kinds
4x=50000
divide by 4
x=12500